# Enhanced microscale NMR spectroscopy of low-gyromagnetic ratio nuclei via hydrogen transfer

### v2

@@ I Introduction
In the liquid state case, NMR spectroscopy targets, basically, two main intramolecular interactions. These are: (i) Chemical shifts, that reveal the chemical environment around certain nucleus in a molecule, acting as fingerprints to identify specific chemical groups in the compound. (ii) J-couplings, which carry information about molecular structure and conformation through bond connectivity, as well as of the angle and distance between nuclei contreras2000angular ; smith1991analysis ; schmidt2010correlation ; schmidt2011one . Measuring these target quantities for a specific isotope requires using that nucleus as the emitter. For example, estimating the J-coupling between hydrogen and a 31P nucleus in a molecule involves tracking oscillations in the NMR signal emitted by hydrogen glenn2018high . In cases where the interaction of interest does not involve nuclei with a large gyromagnetic ratio (such as hydrogen or fluorine), the experiment’s sensitivity decreases. In addition, sensitivity is further compromised by the low natural abundance of the target nuclei. For example, 13C is significantly less common than 1H, leading to lower amplitudes in 13C NMR compared to 1H NMR balci2005basic . In this direction, operating at high magnetic fields offers several advantages. Under such conditions, the thermal polarization of nuclei increases, enhancing the amplitude of the emitted NMR signal while molecular spectra are simplified munuera2023high ; alsina2024jcoupling .
@@
In microscale NMR, nitrogen vacancy (NV) center ensembles have gained significant attention for their ability to probe low-volume samples, offering potential applications in fields such as high-throughput chemistry mennen2019the . In this scenario, various protocols including quantum heterodyne measurements with large frequency resolution (up to $\sim 1$ Hz) have been reported glenn2018high ; bucher2020hyperpolarization ; arunkumar2021micron ; bruckmaier2023imaging . In addition, recently introduced protocols have garnered attention for their ability to operate at high magnetic fields on the microscale via NV ensembles munuera2023high ; daly2024nutation ; alsina2024jcoupling ; munuera2024high . This is achieved through an external RF drive applied to the target sample, which induces rotations over nuclei and causes the emission of a slowly oscillating NMR signal which can be tracked by the NV ensemble. The frequency such NMR signal is constrained by the available RF power as well as by the nature of the targeted nuclei. Consequently, studying nuclei other than hydrogen or fluorine (which have lower gyromagnetic ratios, resulting into slower frequencies) requires longer detection times. This leads to a significant reduction in sensitivity as a result of sensor decoherence.
@@
In this work, we present a pulse-sequence for performing NMR spectroscopy with NV ensembles over low-gyromagnetic-ratio nuclei by using hydrogens as enhancers of the NMR signal. This approach takes advantage of their high abundance in most molecules and their large gyromagnetic ratio, which leads to stronger coupling with NVs and importantly, to fast emission, with a rate that matches NV coherence times. Additionally, the potential for hydrogen hyperpolarization further motivates this approach bucher2020hyperpolarization ; arunkumar2021micron . The method is suited for the context of microscale samples analyzed with NV ensembles at high magnetic fields, finally resulting in increased sensitivities. The protocol can be employed to detect J-couplings, with a resolution limited by the sample’s $T_{2}$, or to detect J-couplings and chemical shifts, where the resolution is instead limited by $T_{2}^{*}$. By detailed numerical simulations, we show that our approach leads to sensitivity improvements of more than one order of magnitude.
@@ II System
describing the effect of RF fields over the sample.
@@
Figure 1: a) Schematic representation of our pulse-sequence. $\pi$-pulses are shown wider than $\pi/2$-pulses. Pulse axes follow the color code in the top legend. Horizontal lines correspond to the radiation scheme over nucleui in the target molecule, namely $S_{1}$, $I_{1}$, $I_{2}$, etc. The main block, repeated $n$ times, is enclosed in brackets (see main text). b) Diagram of the input and output density matrices for each block repetition. c) (TOP) Standard approach with direct interrogation of nucleus $I_{1}$ and (BOTTOM) corresponding input/output nuclear state for each block.
@@ III Protocol
At the detection stage 3, a number $M$ of $2\pi$ and $-2\pi$ rotations along the $y$-axis are applied to the hydrogen nucleus, each with a duration $t_{\rm RF}$. Due to this rotations, the hydrogen emits an NMR signal, $B_{z}(t)$, which is given by meriles2010imaging
@@ IV Sensitivity
We compare the performance of our protocol with respect to a standard approach where the $I_{1}$ nucleus is directly interrogated, see Fig. 1 c). One can check that, at the $k$-th detection stage of the standard approach, we have (see Appendix C)
@@ VI Conclusions
In summary, we present a protocol that enables the detection of NMR signals encoding parameters of low-gyromagnetic-ratio nuclei (such as 13C) through hydrogen nuclei. By leveraging the high gyromagnetic ratio and the common abundance of hydrogen atoms in target molecules, our protocol significantly enhances the signal-to-noise ratio. It achieves more than one order of magnitude improvement for 13C, while even greater gains are expected for nuclei like nitrogen, which have lower gyromagnetic ratios.
@@ Appendix A Analytical calculation of the dynamics
In the encoding stage we have
@@
In the encoding stage 1 we apply $U_{2}$. After such evolution, the state reads
@@ Appendix B From the sample to the NV
Here $\hbar=1.054\cdot 10^{-34}\>\rm{J\cdot s}$ is the reduced Planck constant, $\gamma_{H}=(2\pi)\times 42.57\>\rm{MHz/T}$ is the gyromagnetic ratio of the hydrogen, $\mu_{0}=4\pi\cdot 10^{-7}\>\rm{H/m}$ the magnetic permeability of free space, $\rho_{H}$ the hydrogen density of the sample, $B_{z}=2\>\rm{T}$ the external magnetic field, $K_{B}=1.38\cdot 10^{-23}\>\rm{J/K}$ the Boltzmann constant, $T=300$ K, and $F_{3}=4.1$ glenn2018high characterizes the geometry of the sample. We take a hydrogen density of $\rho_{H}=6.6\cdot 10^{28}\>m^{-3}$.
@@ Appendix C Standard approach
The fairest comparison between both methods would be to set the encoding periods and number of measurements to have the same duration, $\tau_{1}=\tau$ and $n_{1}=n$. With such assumption, the sensitivity ratio between both methods is

### v3

@@ I Introduction
In the liquid state case, NMR spectroscopy targets, basically, two main intramolecular interactions. These are: (i) Chemical shifts, that reveal the chemical environment around certain nucleus in a molecule, acting as fingerprints to identify specific chemical groups in the compound. (ii) J-couplings, which carry information about molecular structure and conformation through bond connectivity, as well as of the angle and distance between nuclei contreras2000angular ; smith1991analysis ; schmidt2010correlation ; schmidt2011one . Measuring these target quantities for a specific isotope requires using that nucleus as the emitter. For example, estimating the J-coupling between hydrogen and a 31P nucleus in a molecule involves tracking oscillations in the NMR signal emitted by hydrogen glenn2018high . In cases where the interaction of interest does not involve nuclei with a large gyromagnetic ratio (such as hydrogen or fluorine), the experiment’s sensitivity decreases. In addition, sensitivity is further compromised by the low natural abundance of the target nuclei. For example, 13C is significantly less common than 1H, leading to lower amplitudes in 13C NMR compared to 1H NMR balci2005basic . n conventional NMR, techniques such as Insensitive Nucleus Enhancement by Polarization Transfer (INEPT) morris1979enhancement are commonly used to transfer the higher thermal polarization of hydrogen nuclei to less sensitive nuclei like 13C, thereby enhancing their polarization and improving detection sensitivity. These methods are also applicable in microscale NMR using NV centers, as they depend only on applying control pulses to the nuclear sample. Another approach is to increase the static magnetic field and operate in a high-field regime. Under such conditions, the thermal polarization of nuclei increases, enhancing the amplitude of the emitted NMR signal while molecular spectra are simplified munuera2023high ; alsina2024jcoupling .
@@
In microscale NMR, nitrogen vacancy (NV) center ensembles have gained significant attention for their ability to probe low-volume samples, offering potential applications in fields such as high-throughput chemistry mennen2019the . In this scenario, various protocols including quantum heterodyne measurements with large frequency resolution (up to $\sim 1$ Hz) have been reported glenn2018high ; bucher2020hyperpolarization ; arunkumar2021micron ; bruckmaier2023imaging . In addition, recently introduced protocols have garnered attention for their ability to operate at high magnetic fields on the microscale via NV ensembles munuera2023high ; daly2024nutation ; alsina2024jcoupling ; munuera2024high . This is achieved through an external RF drive applied to the target sample, which induces rotations over nuclei and causes the emission of a slowly oscillating NMR signal which can be tracked by the NV ensemble. The frequency such NMR signal is constrained by the available RF power as well as by the nature of the targeted nuclei. Consequently, studying nuclei other than hydrogen or fluorine (which have lower gyromagnetic ratios, resulting into slower Rabi frequencies of the applied RF) requires longer detection times. This leads to a significant reduction in sensitivity as a result of sensor decoherence.
@@
In this work, we present a pulse-sequence for performing NMR spectroscopy with NV ensembles over low-gyromagnetic-ratio nuclei by using hydrogens as enhancers of the NMR signal. This approach exploits their high abundance in most molecules and large gyromagnetic ratio, which leads to higher initial polarization and higher coupling with the NV -resulting in fast emission, at a rate matching NV coherence times. Thus, our method not only leverages the high initial polarization of hydrogen nuclei, as in INEPT-like techniques in standard NMR, but also exploits their stronger coupling with the NV ensemble compared to other nuclei. This results in a faster and more intense emitted NMR signal, reducing the detrimental effects of NV decoherence during the measurement period. Additionally, the typically higher concentration of hydrogen relative to other species further enhances the amplitude of the emitted NMR signal. Moreover, the potential for hydrogen hyperpolarization further motivates this approach bucher2020hyperpolarization ; arunkumar2021micron . The method is suited for the context of microscale samples analyzed with NV ensembles at high magnetic fields, finally resulting in increased sensitivities. The protocol can be employed to detect J-couplings, with a resolution limited by the sample’s $T_{2}$, or to detect J-couplings and chemical shifts, where the resolution is instead limited by $T_{2}^{*}$. Through detailed numerical simulations, we demonstrate that our approach improves sensitivity by more than an order of magnitude compared to the only existing standard methods in the scenario of high magnetic fields munuera2023high ; alsina2024jcoupling .
@@ II System
describing the effect of RF fields over the sample. Note that Eq. (1) does not include any term corresponding to the NV sensor, as it only plays a role during the detection of the sample’s evolution (see the next section). Additionally, dipole-dipole and other orientation-dependent interactions are neglected, as they average out in liquids.
@@
Figure 1: a) Schematic representation of our pulse-sequence. The pulse width corresponds to the applied rotation angle (i.e. $\pi/2$ and $\pi$ pulses, as well as the $2\pi$ rotations applied during the detection stages). The $2\pi$ rotations happen only during the detection stages ($\Omega_{H}\cdot t_{\rm RF}=2\cdot 2\pi$) and are not shown to scale relative to the $\pi/2$ and $\pi$ pulses to avoid excessive spacing in the scheme. Pulse axes follow the color code in the top legend. Horizontal lines correspond to the radiation scheme over nucleui in the target molecule, namely $S_{1}$, $I_{1}$, $I_{2}$, etc. The main block, repeated $n$ times, is enclosed in brackets (see main text). b) Diagram of the input and output density matrices for each block repetition. c) (TOP) Standard approach with direct interrogation of nucleus $I_{1}$ and (BOTTOM) corresponding input/output nuclear state for each block.
@@ III Protocol
At the detection stage 3, a number $M$ of $2\pi$ and $-2\pi$ rotations along the $y$-axis are applied to the hydrogen nucleus, each with a duration $t_{\rm RF}$. Due to this rotations and to the scenario we consider -NV ensemble at a sufficient distance from the sample- the interaction between the sample and the NV ensemble (dipole-dipole) can be can be described by a classical NMR signal emitted from the sample $B_{z}(t)$ meriles2010imaging , with the form
@@ IV Sensitivity
We compare the performance of our protocol with respect to a standard approaches munuera2023high ; alsina2024jcoupling where the $I_{1}$ nucleus is directly interrogated, see Fig. 1 c). Note that the experimental setup is the same in all cases: a microscale sample positioned above an NV ensemble that detects its emitted magnetic field. The only thing that changes is the control RF applied to the sample.
One can check that, at the $k$-th detection stage of the standard approach, we have (see Appendix C)
@@
To finish, note that the sensitivity improvement comes at the cost of reduced resolution, as our method shortens the effective decoherence times of the nuclei in the sample (see Appendix D for explicit expressions).
@@ V Numerical Results
Regarding the sensor, we assume an NV ensemble at $1\mu$m depth, leading to a detectable sample in the order of pL glenn2018high ; meriles2010imaging . The NV concentration is approximately $2\cdot 10^{17}{\rm cm}^{-3}$, with a total active NV volume of $V=\pi\times\left(5\cdot 10^{-5}\right)^{2}\times 13\cdot 10^{-6}=10^{-15}{\rm m}^{-3}$, all extracted from glenn2018high . This results in an estimated total of $\sim 10^{8}$ NV centers.
@@ VI Conclusions
We present a protocol that enables the detection of NMR signals encoding parameters of low-gyromagnetic-ratio nuclei (such as 13C) through hydrogen nuclei. By leveraging the high gyromagnetic ratio and the common abundance of hydrogen atoms in target molecules, our protocol significantly enhances the signal-to-noise ratio. It achieves more than one order of magnitude improvement for 13C, while even greater gains are expected for nuclei with lower gyromagnetic ratios.
@@ Appendix A Analytical calculation of the dynamics
In the loading stage we have
@@
In the loading stage 1 we apply $U_{2}$. After such evolution, the state reads
@@ Appendix B From the sample to the NV
Here $\hbar=1.054\cdot 10^{-34}\>\rm{J\cdot s}$ is the reduced Planck constant, $\gamma_{H}=(2\pi)\times 42.57\>\rm{MHz/T}$ is the gyromagnetic ratio of the hydrogen, $\mu_{0}=4\pi\cdot 10^{-7}\>\rm{H/m}$ the magnetic permeability of free space, $\rho_{H}$ the hydrogen density of the sample, $B_{z}=2\>\rm{T}$ the external magnetic field, $K_{B}=1.38\cdot 10^{-23}\>\rm{J/K}$ the Boltzmann constant, $T=300$ K, and $F_{3}$ characterizes the geometry and position of the sample relative to the NV ensemble meriles2010imaging . The maximum value, $F_{3}=4.1$, is achieved when the diamond is cut so that the [111] direction is perpendicular to the surface. We take a hydrogen density of $\rho_{H}=6.6\cdot 10^{28}\>m^{-3}$.
@@ Appendix C Standard approach
The fairest comparison between both methods would be to set the loading periods and number of measurements to have the same duration, $\tau_{1}=\tau$ and $n_{1}=n$. With such assumption, the sensitivity ratio between both methods is